# Rethinking Robustness in Machine Learning: A Posterior Agreement Approach

### v1

@@ Abstract
The robustness of algorithms against covariate shifts is a fundamental problem with critical implications for the deployment of machine learning algorithms in the real world. Current evaluation methods predominantly match the robustness definition to that of standard generalization, relying on standard metrics like accuracy-based scores, which, while designed for performance assessment, lack a theoretical foundation encompassing their application in estimating robustness to distribution shifts.
In this work, we set the desiderata for a robustness metric, and we propose a novel principled framework for the robustness assessment problem that directly follows the Posterior Agreement (PA) theory of model validation. Specifically, we extend the PA framework to the covariate shift setting by proposing a PA metric for robustness evaluation in supervised classification tasks. We assess the soundness of our metric in controlled environments and through an empirical robustness analysis in two different covariate shift scenarios: adversarial learning and domain generalization. We illustrate the suitability of PA by evaluating several models under different nature and magnitudes of shift, and proportion of affected observations. The results show that the PA metric provides a sensible and consistent analysis of the vulnerabilities in learning algorithms, even in the presence of few perturbed observations.
@@ 1 Introduction
Real-world data analysis problems are often formulated as (potentially intractable) combinatorial optimization tasks, e.g., inferring data clusterings, image segmentations, orderings, embeddings, or parameter estimation of dynamical systems, etc. The stochastic nature of the input data and the computational complexity require resorting to approximated, probabilistic estimates. However, such solutions often suffer from instability in generalization to new observations that contain the same signal but systematically different noise perturbations, a phenomenon often referred to as covariate shift (Quiñonero-Candela et al., 2008). Covariate shift has gained increasing relevance, especially with the rise of deep learning and its astonishing improvements for a wide variety of predictive tasks. The focus of the research community is progressively “shifting” towards modelings covering different levels of heterogeneity in the data. Consequently, new experimental settings have been implemented in order to test the robustness of machine learning models under nontrivial perturbations of the signal.
@@
  • adversarial (intentional) shift: the data is crafted ad-hoc by an adversary with vicious intentions, to specifically hinder the output quality of the algorithm (Carlini & Wagner, 2017b; Biggio & Roli, 2018).
  • out-of-distribution (unintentional) shift: the data is subject to different initial conditions during its measurement (e.g., lighting conditions, orientation, and so on) [^1] In some cases, out-of-distribution is used to define a shift in the target set (Fang et al., 2022). (Koh et al., 2021; Wang et al., 2023a).
Under these settings, the usual standard robustness evaluation procedures consist of comparing the accuracy performance of the model under increasing levels of shift (Cinà et al., 2025; Koh et al., 2021). This procedure is essentially the same as that commonly used to evaluate the generalization capabilities of models in the absence of (measurable/accountable) shift, that is, when data is subject to the randomness entailed by the sampling process only. In the following, we argue that robustness assessment requires a paradigm shift in the way it is approached.
Figure 1: Evolution of accuracy and our proposed metric, Posterior Agreement (PA), displayed by random, constant, and perfect classifiers under a Bernoullian sampling of size $1000$. The random classifier sample is generated by permuting the original so that the number of mismatched observations depends on the Bernoulli probability. Accuracy does not comply with the desired properties of a robustness metric and provides an inconsistent assessment that is exclusively driven by task performance, while PA discriminates the robust from the unrobust classifier, in particular, detecting the robustness of a constant classifier.
First of all, to better characterize the concept of robustness metric, we list two properties that it should possess:
@@
    non-increasing: the metric should be non-increasing with respect to the response of the model against increasing levels of the shift power.
@@
    shift-sensitive: the metric should differentiate models only by their generalization capabilities against covariate shift. For instance, the metric should be independent of the task performance of the model.
@@
The properties described above serve as guiding principles in the robustness evaluation process. However, robustness itself may be susceptible to different interpretations. For example, the concept of shift power can be encoded in several ways. In adversarial learning (Madry et al., 2018), it usually coincides with the norm of the difference between a data observation and its perturbed counterpart. On the other hand, in domain generalization (Eulig et al., 2021), shifts usually encode perceptual or semantic alterations that are specific to each dataset and that are hard or even impossible to characterize in a unified way. Furthermore, the shift power may also describe different aspects of the perturbation, such as the number of affected observations, often neglected but equally relevant to time and performance efficiency. In both settings, a principled approach to measuring robustness is required and is currently missing.
@@
In general, Property 1 is satisfied by accuracy-based metrics. However, as accuracy results from thresholding over the predictions, it is unable to encapsulate any information regarding model confidence. In standard validation procedures, devising prediction-confidence-based metrics increases their discriminative power, for instance, when comparing models with similar predictive capabilities but different confidence. In the covariate shift setting, this distinction is decisive, not only for detecting low shift power perturbations but also for providing a more consistent robustness assessment that does not heavily depends on the stochasticity of the dataset.
Figure 2: The optimization over the inverse temperature parameter $\beta$ is required to make the two posteriors insensitive to noise perturbations. In the top figures, the curves represent two posteriors $p(c\mid X^{\prime})$, $p(c\mid X^{\prime\!\>\!\prime})$ in this case over a real-valued parameter $c$, (e.g., a distribution’s mean). When $\beta$ is too low (top left), the posteriors tend to be uniform and uninformative. When the $\beta$ is too high (top right) the posteriors are too peaked and therefore sensitive to even small noise perturbation. The optimal $\beta$ (top center) maximizes the posterior agreement (bottom) guaranteeing informativeness and stability in the distributions that can be then used for model selection or, in our case, for robustness assessment.
Additionally, some accuracy metrics involve a comparison with the target variables in the test set (e.g., the attack success rate in adversarial learning), thus not complying with Property 2. For example, a constant classifier outputting the same prediction, regardless of the input, is robust by definition, since its output is independent of any shift in the data. Similarly, a perfect classifier that always outputs the correct prediction for any input is also robust. An accuracy metric would, however, discriminate the two models as differently robust since their performance would differ (cf. Figure 1 left).
@@
In this work, we overcome these weaknesses and propose a new metric complying with the aforementioned properties, from the principled perspective of the Posterior Agreement (PA) framework (Buhmann, 2010; Buhmann et al., 2018; Gronskiy, 2018). PA is an alternative model validation procedure rooted in information-theoretic thinking, in particular the rate-distortion theory (Cover & Thomas, 1999). In this setting, the learning algorithm is thought of as a lossy compression procedure with the aim of resolving the hypothesis class as precisely as possible, given the stochastic data source, a critical requirement for a metric aligned with the two proposed properties.
The proposed PA metric provides a unique and unified framework for robustness assessment in the covariate shift setting, as it relies on a concept of robustness that does not stem from the nature of the data or the shift but instead from the consistency of the probabilistic response of the model. Therefore, PA is a confidence-based metric that does not depend on a model’s classification performance, thus aligning with the foundational properties outlined before. As an illustrative example, Figure 1 depicts the difference between accuracy and PA in the robustness assessment of the constant/perfect classifiers described before.
Following epistemologically grounded thinking, we assume two datasets $X^{\prime}$, ${X}^{\prime\!\>\!\prime}\in\mathcal{X}$, derived from the same experiment, both with ideally identical signal, but different noise realizations. Given a hypothesis set $\mathcal{C}$, the posterior distributions $p(c\mid X^{\prime})$, $p(c\mid X^{\prime\!\>\!\prime})$, for $c\in\mathcal{C}$, should become as similar as possible when adapting the machine learning algorithms and, as a consequence, the functional form of the posterior $p(\cdot\mid\cdot)$ is optimized for succinctness and stability. Traditionally, posteriors have been designed by defining a cost function and then by calculating or approximating the respective Gibbs distribution. The spread of such a posterior is controlled by a width parameter $\beta$, also known as inverse computational temperature. An optimal control of $\beta$ should render the posterior as much as possible insensitive to noise perturbation (cf. Figure 2).
@@
In this article, we first adapt the PA framework to be applied to covariate shift settings without modifying its fundamental principles. In particular, we formulate PA without any requirements of supervised information, making it a suitable metric for robustness evaluation in contexts where the supervision is very scarce or null (e.g., medical applications, or autonomous driving). We then conduct a robustness analysis on two current settings of interest, namely adversarial learning in the evasion attack scenario and domain shift under targeted covariate shifts. In particular, we analyze the performance of several models in terms of PA under different nature and magnitudes of perturbation and proportion of affected observations. Not only does PA succeed in documenting the worsening of model performance in great detail, but it also detects an ongoing shift even when the number of targeted observations is low. Additionally, these findings were extended by leveraging PA for robust model selection with early stopping. In particular, experiments included the perturbation of data with spurious factor co-occurrences. The results show that PA-based selection succeeds at mitigating the spurious correlations existing between shifted environments.
@@
In summary, (i) we propose the PA metric, to assess model robustness in covariate shift scenarios and motivate it theoretically, (ii) we conduct a robustness analysis on several settings, illustrating the findings emerging from an evaluation with PA, (iii) we summarize our findings and discuss a new notion of model robustness in the PA sense arising from the presented results.
@@ 2 Related Work
Adversarial robustness is a fundamental research area in machine learning due to the susceptibility of deep neural networks to small perturbations, leading to erroneous predictions (Szegedy et al., 2014; Goodfellow et al., 2015). A huge plethora of effective and efficient attacks such as FGSM (Goodfellow et al., 2015; Yuan et al., 2019), PGD (Kurakin et al., 2018; Madry et al., 2018), BIM (Kurakin et al., 2018), C&W (Carlini & Wagner, 2017b), DeepFool (Moosavi-Dezfooli et al., 2016), FMN (Pintor et al., 2021), and others (Carlini & Wagner, 2017a; Croce & Hein, 2019; Modas et al., 2019; Zheng et al., 2023), have been developed to evaluate the adversarial robustness of machine learning models. Adversarial examples are often measured in terms of their distance from the original input, using $\ell_{p}$ norms (Carlini & Wagner, 2017b). Their use allows for a consistent and well-defined way to measure the magnitude of the perturbation and to compare different adversarial attacks (Carlini & Wagner, 2017b; Cinà et al., 2025). In this work, we also include analyses on the ratio of perturbed observations in a dataset. We believe that this aspect can help in better understanding the behavior of models against attacks, as later shown in the experimental section.
Assessing model robustness proved to be necessary for developing more reliable machine learning systems and for ensuring their resilience before deployment. Typically, white-box models are employed to analyze a model’s worst-case scenarios and avoid relying on security by obscurity (Carlini et al., 2019; Eisenhofer et al., 2021; Däubener et al., 2020). The assessment is commonly performed in the literature by scoring models with accuracy-based metrics such as attack success rate and adversarial accuracy. While the definition of these two scores is not consolidated, in general, they estimate the model performance over a test set in terms of un/successful attacks. Additionally, Weng et al. (2018) approach robustness evaluation as a Lipschitz constant estimation problem. The method is, however, susceptible to gradient masking (Goodfellow, 2018), causing the metric to overestimate the size of the perturbation needed to fool a model. In addition, the bound is obtained by exploiting the ReLU property of neural networks, therefore narrowing its scope to such models. Wang et al. (2023b) uses, instead, the converging time to an adversarial sample for estimating the robustness of a model. This approach requires the estimation of a Jacobian with respect to the input, which may be infeasible in the case of large-sized data, and is applicable only to differentiable models.
@@
Domain generalization is a field of paramount importance in machine learning due to both the limitation of acquiring a diverse and large enough set of training samples (Wolpert & Macready, 1997), and the nature of current methods to capture spurious correlations from training data that lead to catastrophic loss of performance in out-of-distribution settings (Wiles et al., 2022; Geirhos et al., 2020). Over the years, several benchmark datasets have been introduced specifically for domain generalization. The PACS dataset (Yu et al., 2022) represents shifts in visual styles between real photos and artistic media like paintings, cartoons, and sketches, reflecting variations in colour, texture, and abstractness. VLCS (Torralba & Efros, 2011) captures object-centric and scene-centric distribution shifts, with variations in object variability, scene context, and annotation styles across other previously established datasets. The WILDS benchmark (Koh et al., 2021) embodies real-world shifts, covering temporal and geographical changes, as well as differences in image resolution and quality, which are typical in practical applications. Each benchmark encapsulates unique distribution shifts, providing diverse challenges for evaluating domain generalization methods. More recently, the DiagViB-$6$ dataset (Eulig et al., 2021) has also been proposed for the systematic analysis of data distribution shifts over multiple generative factors, covering hue, position, lightness, scale, and texture, and allowing for accumulation of multiple shifts. While several new benchmarks have been proposed to address growing requirements for both more realistic and challenging evaluation settings, accuracy on the unseen domains remains the fundamental metric for estimating the model robustness.

### v2

@@ Abstract
The robustness of algorithms against covariate shifts is a fundamental problem with critical implications for the deployment of machine learning algorithms in the real world. Current evaluation methods predominantly measure robustness through the lens of standard generalization, relying on task performance measures like accuracy. This approach lacks a theoretical justification and underscores the need for a principled foundation of robustness assessment under distribution shifts. In this work, we set the desiderata for a robustness measure, and we propose a novel principled framework for the robustness assessment problem that directly follows the Posterior Agreement (PA) theory of model validation. Specifically, we extend the PA framework to the covariate shift setting and propose a measure for robustness evaluation. We assess the soundness of our measure in controlled environments and through an empirical robustness analysis in two different covariate shift scenarios: adversarial learning and domain generalization. We illustrate the suitability of PA by evaluating several models under different nature and magnitudes of shift, and proportion of affected observations. The results show that PA offers a reliable analysis of the vulnerabilities in learning algorithms across different shift conditions and provides higher discriminability than accuracy-based measures, while requiring no supervision.
@@ 1 Introduction
Real-world data analysis problems are often formulated as (potentially intractable) combinatorial optimization tasks, e.g., inferring data clusterings, segmentations, embeddings, parameter estimation, etc. The stochastic nature of the input data and the computational complexity of the involved information processing require resorting to approximated, probabilistic estimates. However, such solutions often suffer from instability when generalizing to new observations that contain the same signal but systematically different noise perturbations, a phenomenon often referred to as covariate shift (Quiñonero-Candela et al., 2008). Covariate shift has gained increasing relevance, especially with the rise of deep learning and its astonishing improvements for a wide variety of predictive tasks. The focus of the research community is progressively “shifting” towards novel models that cover different levels of heterogeneity in the data and their structure. Consequently, new experimental settings have been implemented in order to test the robustness of machine learning models under nontrivial perturbations of the signal.
@@
  • adversarial (intentional) shift: data is crafted ad-hoc by an adversary with vicious intentions, to hinder the output quality of the algorithm (Carlini & Wagner, 2017b; Biggio & Roli, 2018).
  • out-of-distribution (unintentional) shift: data is subject to different initial conditions during its measurement (e.g., lighting conditions, orientation, and so on) [^1] In some cases, out-of-distribution is used to define a shift in the target set (Fang et al., 2022). (Koh et al., 2021; Wang et al., 2023a).
Under these settings, the usual standard robustness evaluation procedures consist of comparing the task performance of the model by computing a performance measure – i.e., the accuracy or a derived alternative – under increasing levels of shift (Cinà et al., 2025; Koh et al., 2021). Therefore, such an evaluation is the same commonly used to test the generalization capabilities of models in the absence of measurable/accountable shift, that is, when data is subject to the randomness entailed by the sampling process only. In the following, we argue that the assessment of model robustness requires a paradigm shift in the way it is approached.
Figure 1: Comparison of accuracy and our proposed measure, Posterior Agreement (PA), to assess the robustness of three binary classifiers. We simulate a dataset of targets $D_{Y}=\{y_{1},\dots,y_{n}\}$, $|D_{Y}|=N$ where each $y_{i}\sim\mathcal{B}(p)$ is sampled from a Bernoulli distribution with $p=p(Y^{\prime}=1)$ (displayed on the $x$-axis). We assume a perfect classifier $f(y_{i})=y_{i}$, a constant classifier $f(y_{i})=0$, and a random classifier whose outputs are a permutation of $D_{Y}$, so that the number of mismatched observations depends on $p$. Accuracy does not comply with the desired properties of a robustness measure and provides an inconsistent assessment, which is exclusively driven by task performance. Instead, PA detects the robustness of a constant classifier, discriminating it from the random, unrobust one.
We first characterize the concept of robustness measure with two ideal properties that it should possess:
@@
    non-increasing: the measure should be non-increasing with respect to the response of the model against increasing levels of the shift power.
@@
    shift-sensitive: the measure should differentiate models only by their response against covariate shift.
In essence, Property 1 requires the measure not to increase the robustness score of a model if the shift in the data increases. In the best case, a model should be insensitive to shift, therefore scoring the same robustness as when it is absent. Property 2 requires, instead, that the measure reacts only to covariate shift and is not, for instance, sensitive to in-distribution sampling randomness and its related task performance.
@@
Such properties serve as guiding principles for the evaluation process and are intentionally not formalized. This qualitative reasoning is due to the concept of robustness depending, case-by-case, on the definition of shift power. For example, in adversarial learning (Madry et al., 2018), the shift power usually coincides with the norm of the difference between a data observation and its perturbed counterpart. On the other hand, in domain generalization (Eulig et al., 2021), it usually encodes perceptual or semantic alterations that are specific to each dataset and that are hard or even impossible to characterize in a unified way. Furthermore, the shift power may also describe different aspects of the perturbation, such as the number of affected observations, which is often overlooked but equally relevant to time and performance efficiency. In both settings, a principled approach to measuring robustness is required and is currently missing. Another difficulty, related to Property 2, arises from the fact that it is not always possible to separate the covariate shift from the sampling noise, as this would require access to the often unknown data-generating process.
@@
In general, accuracy-based measures decline with growing shift power, thereby satisfying Property 1. However, as accuracy results from thresholding over the predictions, it is unable to encapsulate any information regarding model confidence. In standard validation procedures, devising prediction-confidence-based measures increases their discriminative power, especially when comparing models with similar predictive capabilities but different confidence in the predictions. In the covariate shift setting, this distinction is decisive, not only for characterizing a model’s robustness even in the presence of low shift power perturbations but also for providing a more consistent robustness assessment that does not heavily depend on the stochasticity of the dataset. Additionally, some accuracy measures involve a comparison with the target variables in the test set (e.g., the attack success rate in adversarial learning), thus not complying with Property 2. For example, a constant classifier outputting the same prediction, regardless of the input, is robust by definition, since its output is independent of any shift in the data, thus covariate shift too. Similarly, a perfect classifier that always outputs the target’s input is also robust. An accuracy measure would, however, discriminate the two models as differently robust since their performance would differ (cf. Figure 1, left).
@@
In this work, we address these weaknesses and propose a new measure that better complies with the aforementioned properties, from the principled perspective of the Posterior Agreement (PA) framework (Buhmann, 2010; Buhmann et al., 2018a; Gronskiy, 2018). PA is an alternative model validation procedure rooted in information-theoretic thinking, in particular the rate-distortion theory (Cover & Thomas, 1999). In this setting, the learning algorithm is thought of as a lossy compression procedure with the aim of resolving the hypothesis class as precisely as possible, given the stochastic data source, a critical requirement for a measure aligned with the two proposed properties. The proposed PA measure provides a unique and unified framework for robustness assessment in the covariate shift setting, as it relies on a concept of robustness that stem from the consistency of the probabilistic response of the model under different data instances. PA is a confidence-based measure that does not depend on a model’s classification performance, thus aligning with the foundational properties outlined before. As an illustrative example, Figure 1 depicts the difference between accuracy and PA in the robustness assessment of the constant/perfect classifiers described before. Since, the prediction confidence of a constant model is unaffected by the data, PA awards its robustness.
The original posterior agreement framework was devised to provide an epistemologically grounded methodology for finding robust solutions to optimization problems. In short, the method is given two dataset instances, used to fit two posterior distributions. The spread of such distributions is governed by a temperature parameter that trades off their informativeness and stability, in order to reach an agreement and reconcile their differences due to the sampling noise in the data. In this article, we focus on leveraging the PA framework for robustness evaluation. In particular, we assume that the shift between the two data realizations is covariate. By using the PA approach, we quantify the agreement between the related posteriors. The obtained measure estimates, thus, the robustness of a tested model as its capability to treat the shift as generalizable sampling noise. In such a case, the posteriors strongly agree, and our PA measure awards the model a good score.
@@
Our work proceeds as follows. We first present the original PA framework and adapt it to covariate shift settings without modifying its fundamental principles. In particular, we formulate PA without any requirements of supervised information, making it a suitable measure for robustness evaluation in contexts where the supervision is very scarce or null (e.g., medical applications, or autonomous driving). We then conduct a robustness analysis on two current settings of interest, namely adversarial learning in the evasion attack scenario and domain shift under targeted covariate shifts. In particular, we analyze the performance of several models in terms of PA under different natures and magnitudes of perturbation and proportion of affected observations. Not only does PA succeed in documenting the worsening of model performance in great detail, but it also detects an ongoing shift even when the number of targeted observations is low. Additionally, these findings were extended by leveraging PA for robust model selection with early stopping. In particular, experiments included the perturbation of data with spurious factor co-occurrences. The results show that PA-based selection succeeds at mitigating the spurious correlations existing between shifted environments.
@@
In summary, (i) we propose the PA measure to assess model robustness in covariate shift scenarios and to motivate it theoretically, (ii) we conduct a robustness analysis on several settings, illustrating the findings emerging from an evaluation with PA, (iii) we summarize our findings and discuss a new notion of model robustness in the PA sense arising from the presented results.
@@ 2 Related Work
Posterior Agreement
Since its inception, the PA framework has been adapted to several different settings and is present in the literature under different variants, reflecting its evolution over time. The earliest version is called Approximate Set Coding (ASC, Buhmann, 2010; 2012), that interprets a model selection problem as an imaginary communication channel entailing a capacity optimization process. At the core of ASC lies the trade-off between informativeness and robustness of the selected solution. ASC works with approximation sets, i.e., discrete subsets of hypotheses spanned by a cost function of the problem. Such sets are controlled by a tunable parameter determining an optimal trade-off between informativeness and stability of the solutions, discussed in more detail later. A variant of the framework goes under the name of similarity approach (Buhmann et al., 2018b), where approximation sets are tuned to maximize the ratio of solutions expected due to the similarity of the data instances. A recent version of the framework (Buhmann et al., 2018a) applies a relaxation of the ASC approach with posterior distributions that will be presented in this work. Here, approximation sets are replaced by Gibbs posteriors tunable through a temperature parameter. In the following, we will refer to this method as PA. In all its variants, PA has been applied to a variety of different selection and validation problems, including clustering (Buhmann, 2010), regression (Gorbach et al., 2017; Wegmayr & Buhmann, 2021), sorting (Busse et al., 2013), MaxCut problems (Gronskiy & Buhmann, 2014; Bian et al., 2016), and assignment problems (Buhmann et al., 2014; 2017; 2018a).
@@
Adversarial robustness is a fundamental research area in machine learning due to the susceptibility of deep neural networks to small perturbations, leading to erroneous predictions (Szegedy et al., 2014; Goodfellow et al., 2015). A huge plethora of effective and efficient attacks such as FGSM (Goodfellow et al., 2015; Yuan et al., 2019), PGD (Kurakin et al., 2018; Madry et al., 2018), BIM (Kurakin et al., 2018), C&W (Carlini & Wagner, 2017b), DeepFool (Moosavi-Dezfooli et al., 2016), FMN (Pintor et al., 2021), and others (Carlini & Wagner, 2017a; Croce & Hein, 2019; Modas et al., 2019; Zheng et al., 2023), have been developed to evaluate the adversarial robustness of machine learning models. Adversarial examples are often measured in terms of their distance from the original input, using $\ell_{p}$ norms (Carlini & Wagner, 2017b). Their use enables a consistent and well-defined way to measure the magnitude of the perturbation and to compare different adversarial attacks (Carlini & Wagner, 2017b; Cinà et al., 2025). In this work, we also include analyses on the ratio of perturbed observations in a dataset. We believe that this aspect can help in better understanding the behavior of models against attacks, as later shown in the experimental section.
Assessing model robustness proved to be necessary for developing more reliable machine learning systems and for ensuring their resilience before deployment. Typically, white-box models are employed to analyze a model’s worst-case scenarios and avoid relying on security by obscurity (Carlini et al., 2019; Eisenhofer et al., 2021; Däubener et al., 2020). The assessment is commonly performed in the literature by scoring models with accuracy-based measures such as attack success rate and adversarial accuracy. While the definition of these two scores is not consolidated, in general, they estimate the model performance over a test set in terms of un/successful attacks. Additionally, Weng et al. (2018) approach robustness evaluation as a Lipschitz constant estimation problem. The method is, however, susceptible to gradient masking (Goodfellow, 2018), causing the measure to overestimate the size of the perturbation needed to fool a model. In addition, the bound is obtained by exploiting the ReLU property of neural networks, therefore narrowing its scope to such models. Wang et al. (2023b) uses, instead, the converging time to an adversarial sample for estimating the robustness of a model. This approach requires the estimation of a Jacobian with respect to the input, which may be infeasible in the case of large-sized data, and is applicable only to differentiable models.
@@
Domain generalization is a field of paramount importance in machine learning due to both the limitation of acquiring a diverse and large enough set of training samples (Wolpert & Macready, 1997), and the nature of current methods to capture spurious correlations from training data that lead to catastrophic loss of performance in out-of-distribution settings (Wiles et al., 2022; Geirhos et al., 2020). Over the years, several benchmark datasets have been introduced specifically for domain generalization. The PACS dataset (Yu et al., 2022) represents shifts in visual styles between real photos and artistic media like paintings, cartoons, and sketches, reflecting variations in colour, texture, and abstractness. VLCS (Torralba & Efros, 2011) captures object-centric and scene-centric distribution shifts, with variations in object variability, scene context, and annotation styles across other previously established datasets. The WILDS benchmark (Koh et al., 2021) embodies real-world shifts, covering temporal and geographical changes, as well as differences in image resolution and quality, which are typical in practical applications. Each benchmark encapsulates unique distribution shifts, providing diverse challenges for evaluating domain generalization methods. More recently, the DiagViB-$6$ dataset (Eulig et al., 2021) has also been suggested for the systematic analysis of data distribution shifts over multiple generative factors, covering hue, position, lightness, scale, and texture, and allowing for the accumulation of multiple shifts. While several new benchmarks have been proposed to address growing requirements for both more realistic and challenging evaluation settings, accuracy on the unseen domains remains the fundamental measure for estimating model robustness.